OpenStudy (anonymous):

given that p and 1 are the roots of the quadratic equation x^2-(2k-3)x+3=0. find the value of p and k. b) given that the quadratix\c equation 4x^2-(8k-12)x-8k^2+6k+21=0 has two different real roots. find the range of values of k.

OpenStudy (anonymous):

It has been a while since i have used these so I don't want to give you a wrong answer, I will try to get someone here that knows how to do it :P

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Maybe u can use following formula for sum and product of roots sum = -b/a product = c/a

OpenStudy (anonymous):

@estudier I was just about to write that ^^

OpenStudy (anonymous):

Sorry, u carry on then...

OpenStudy (anonymous):

... lol It seems that I can't do it, I got stuck in a minus I'm deeply in doubt what to do ! @_@

OpenStudy (anonymous):

\[\LARGE x^2+px+k=0\] \[\LARGE x_1+x_2=-p\] \[\LARGE x_1\cdot x_2=k\] \[\LARGE x^2-(2k-3)x+3=0\] \[\LARGE x_1+x_2=-p\] \[\LARGE x_1+x_2=-(2k-3)\] \[\LARGE -p=-(2k-3)\] \[\LARGE p=2k-3\] \[\LARGE x_1\cdot x_2=k\] \[\LARGE x_1\cdot x_2=3\] \[\LARGE k=3\] \[\LARGE p=2k-3\] \[\LARGE p=2*3-3\] \[\LARGE p=3\] so p=k could it be?? ( I'M NOT SURE ABOUT THIS SOLUTION) I have to go, see you later ...